Problem 16 that i got wrong
Find the slope of any line perpendicular to the line through points (0,5) and (-3, -4)

Answer by venugopalramana(3286)

Problem 16 that i got wrong
Find the slope of any line perpendicular to the line through points (0,5)...(X1,Y1) and (-3, -4)..(X2,Y2)
SLOPE OF THIS LINE =(Y2-Y1)/(X2-X1)=(-4-5)/(-3-0)=3
HENCE SLOPE OF PERPENDICULAR=-1/3